OpenStudy (anonymous):

use natural log to solve 3^2x=20

OpenStudy (anonymous):

Do you know logarithmic properties? http://www.andrews.edu/~calkins/math/webtexts/numb17.htm

OpenStudy (anonymous):

sort of! I tried using a bunch of different websites to help but I have no idea how to approach the problem

OpenStudy (anonymous):

\[3^{2x}=20\]\[\log~3^{2x}=\log~20\]\[2x~\log~3=\log~20\]\[2x=\log_3~20\]\[x= \frac{1}{2}\log_3~20\]\[x=\log_3~20^{1/2}\]\[x=\log_3~\sqrt{20}\]\[x=\log_3~2\sqrt{5}\]\[x=\log_3~2+\frac{1}{2}\log_35\]

OpenStudy (anonymous):

its a multiple choice problem, the answers given are a. ln20/ln6 b. ln(10/3) c. ln20/ln9 or d. ln20-ln3/2